A bowl of water has a temperature of 50◦C. It is put into a refrigerator where the temperature is 5◦C. After 0.5 hours, the water is stirred and its temperature is measured to be 20◦C. It is then left to cool further, take t in Newton’s law of cooling to be measured in
minutes. Use Newton’s law of cooling to predict when the temperature will be 10◦C.
1
Expert's answer
2020-07-01T19:03:57-0400

Solution:

T(t)= Ts+(T0-Ts)e-kt

T(t) =temperature of the object at time t

Ts= temperature of surrounding

T0=temperature of the object at time

 t=0

k=constant


A bowl of water has a temperature of 50◦C. It is put into a refrigerator where the

 temperature is 5◦C.After 0.5 hours, 

the water is stirred and its temperature

is measured to be 20◦C. So we have:


20= 5+(50-5)e-0.5k


e-0.5k="\frac{1}{3}"


-(0.5)k= In"\frac{1}{3}"


k= In 9

It is then left to cool further. 

We should predict when the temperature will be 10◦C:


10= 5+(20-5)e-(In 9)t


e-(In 9)t ="\frac{1}{3}"

t= 0.5

The temperature will be 10◦C in 30 minutes

(after we put a bowl of water into a

 refrigerator the second time).